Researchers have created every possible "knockout" line in yeast. Each line has exactly one gene deleted and all the other genes present (Steinmetz et al. 2002). The growth rate - how fast the number of cells increases per hour - of each of these yeast lines has also been measured, expressed as a multiple of the growth rate of the wild type that has all the genes present. In other words, a growth rate greater than 1 means that a given knockout line grows faster than the wild type, whereas a growth rate less than 1 means it grows more slowly. Below is the growth rate of a random sample of knockout lines:

0.79, 0.88, 0.61, 0.65, 0.86, 0.68, 0.72, 0.91, 0.97, 0.74, 0.96

What is the variance of growth rate this sample of yeast lines (answer to 6 decimals)?
Answer #1

From the given data

S.NO. Sample 1 (X) X^2
1 0.79 0.6241
2 0.88 0.7744
3 0.61 0.3721
4 0.65 0.4225
5 0.86 0.7396
6 0.68 0.4624
7 0.72 0.5184
8 0.91 0.8281
9 0.97 0.9409
10 0.74 0.5476
11 0.96 0.9216
Total: 8.77 7.1517

X--,-8.77 0.797273 n 11 anance(s2)-|z?-n-21=1[7.1517-11*0.7972732]=0015962 SD(s)-V0.0 15962 = 0.12634 2-n-2 --Γ7. 15 17-11 * 0.7972732-0015962